OpenStudy (anonymous):

Solve 4x2 = 12x + 40.

OpenStudy (katrinakaif):

Bring it all onto one side: 4x^2 -12x -40 = 0 Take out the GCF Then Factor

OpenStudy (anonymous):

i believe the answer ur looking for is this: Divide both sides by 4: x² - 3x = 10 Subtract 10 from both sides to get everything on the left: x² - 3x - 10 = 0 Factor: (x - 5)(x + 2) = 0 Answer: x = 5 or x = -2